OpenStudy (anonymous):

you split 2000 between two accounts, account a pays 2% annual interest and account b pays 5% annual interest after one year you have earned a total of 77.50 in interest how much did you invest in each account. i have been stuck on this for hours, can someone help me please? will give medal and fan if i get help.

OpenStudy (anonymous):

@blondie$$$

OpenStudy (anonymous):

r u just trying to figure out how much you'd have

OpenStudy (anonymous):

im trying to find out how much money is in each account

OpenStudy (anonymous):

oh ok..was there a formula to go with it?

OpenStudy (anonymous):

if possible it would be awesome for me to see the work so that i know how to do it next time.

OpenStudy (anonymous):

yeah kind of, x+y=2000 and the other one is 2/100+5/100=77.50

OpenStudy (anonymous):

can u just fill in x+y with the ammount of money?? or divide

OpenStudy (anonymous):

iv tried that, theres a way to do it, i had a problem like this before and my teacher walked me through it.

OpenStudy (anonymous):

oh ok...@dan815

OpenStudy (anonymous):

im trying to think about what i can do

OpenStudy (anonymous):

got it, if you multiply the lowest % in this case its 2% you get 2X+2Y=4000, subtract that from 2X+5Y=7750 than you get 3Y=3750, devide that by 3 and you get Y=1250 so X must be 750. so the answer is Y=1250, X=750.

OpenStudy (anonymous):

can someone check my work to make sure i didnt mess up any calculations?

OpenStudy (anonymous):

yeah just a sec xD... @dan815

OpenStudy (campbell_st):

well the simple solution is let x be the amount earning 2% then the amount earning 5% is 2000 - x remember the percentages need to be expressed as decimals and their totals is the interest 0.02x + 0.05(2000 - x) = 77.50 which means 0.02x + 100 - 0.05x = 77.50 or 100 - 0.03x = 77.50 now solve for x... remember this is the amount at 2% then subtract this amount from $2000 to find the investment at 5% hope it makes sense